# G-Transformer: Counterfactual Outcome Prediction under Dynamic and Time-varying Treatment Regimes

### v2

@@ 5.3 Predictive Check: Population-Level and Individual-Level RMSE
We used G-Transformer to simulate covariate trajectories for patients in the test dataset under two conservative fluids strategies (with 3L and 5L fluid cap) and a fluid liberal strategy (with no cap). Selected covariates are presented in Figure 5, allowing for comparison of the three counterfactual regimes. Population-level average trajectories were calculated from the test set (N=894). Counterfactual strategies are applied starting at time step 1, conditioned on observations up until the first hour in the ICU. For each patient, we perform 100 Monte Carlo simulations. The plotted values are averaged across simulated trajectories for all test patients.
Our aim is to assess whether G-Transformer can make logical and clinically meaningful counterfactual predictions. The trends depicted in Figure 5 align with our physiological expectations. Under fluid liberal strategies, the bolus volume, urine output, respiratory rate, and blood pressure exhibit higher values, following the expected trends at a population level. Conversely, the higher levels in creatinine and blood urea nitrogen (BUN) noted under the fluids conservative regime also align with expected outcomes.
Readers should exercise caution when interpreting these plots in Figure 5 as indicators of treatment effects, as the number of patients (or trajectories) contributing to the average may not remain constant throughout the entire 24-hour period. Simulated trajectories might terminate prematurely within the 24-hour period due to death or discharge.
@@
Figure 5: Population-level trajectories under counterfactual fluids strategies. G-Transformer’s predictions for selected covariates under various counterfactual fluids strategies, including two fluids limiting strategies with 3L (blue) and 5L (purple) fluids cap, and a fluids liberal (green) strategy without fluid cap.
@@ 6 Discussion and Conclusion
In this paper, we present G-Transformer, a Transformer-based framework supporting g-computation, to estimate clinical outcomes under counterfactual treatment strategies from both synthetic dataset and a real-world sepsis patient cohort. Our experiments using Cancer Growth and CVSim data, demonstrated that G-Transformer achieved superior performance over other state-of-the art approaches. Using a real-world sepsis dataset, we demonstrated that G-Transformer provides clinically meaningful forecasts of covariate trajectories under alternative counterfactual fluid limiting regimes. This study highlights G-Transformer’s potential clinical utility in supporting treatment decisions for ICU patients. Although we focus on applications in healthcare in this study, the G-Transformer architecture can be applied to other sequential decision making tasks that involve dynamic and time-varying interventions.
@@ A.1 Alternative G-Transformer Architecture Experiment
Our proposed G-Transformer differs from the traditional Transformer, including the Causal Transformer, in the way it is iteratively trained on time-series data. We concatenated the results of the last time step of each sub-sequence to form a complete prediction sequence for optimization. In contrast, traditional Transformers directly use a sequence-to-sequence framework, shifting the training sequence to obtain the sequence that needs to be predicted. We compared the performance of models implemented with these two training methods on the CVSim dataset, and the results are shown in Table 5.
@@
A.2 Gaussian Noise Simulation Experiment
In our proposed G-Transformer Monte Carlo simulation, we sample noise from a distribution of empirical loss from a held-out dataset. Here, we evaluate an alternative approach to sample noise from a normal distribution, with zero mean and empirical covariance estimated from data. As shown in Table 6, sampling noise empirically results in consistently better performance in term of individual-level RMSE across various models when applied to CVSim data under gc1.
Table 6: CVSim: Counterfactual prediction under gc1. Performance of various models with empirical and Gaussian noise sampling in predicting remaining 32-hour patient covariate trajectories based on first 34-hour patient covariate trajectories. Values reported represent the individual-level RMSE between the predicted and counterfactual trajectories.
Empirical Noise Gaussian Noise
Model Individual-level RMSEgc1 Individual-level RMSEgc1
Linear(g-comp) 1.103 1.418
G-Net 1.020 1.158
G-Transformer 1.016 1.240
A.3 Individual-Level RMSE Over Time
@@
Appendix B Additional MIMIC Experiments and Results
@@ Appendix C CVSim Data Generation
A CVSim 6-compartment circulatory model takes as inputs 28 variables that together govern a hemodynamic system. It then deterministically simulates forward in time a set of 25 output variables according to a collection of differential equations (parameterized by the input variables) modeling hemodynamics. Important variables in CVSim include arterial pressure (AP), central venous pressure (CVP), total blood volume (TBV), and total peripheral resistance (TPR). In real patients, physicians observe AP and CVP and seek to keep them above a clinically safe threshold. They do this by intervening on TBV (through fluid administration) and TPR (through vasopressors).
We defined simulated treatment interventions that were designed to mimic the impact of fluids and vasopressors. These simulated interventions alter the natural course of the simulation by increasing either TBV (in the case of the simulated fluids intervention) or TPR (in the case of the simulated vasopressor intervention). We generated patients by randomly initiating baseline inputs (which we hid from our G-Nets to make this a stochastic modeling problem) within plausible physiologic ranges, then using CVSim to simulate covariates forward, intervening according to the relevant treatment strategy at each timestep. Full details of the simulation process can be found in the Appendix.
@@
Appendix D MIMIC Data and Experiments
@@
k Edema MV Diuretics Dialysis Release Death
2 (894, 29.98%) (894, 40.94%) (894, 15.10%) (894, 2.68%) (894, 13.20%) (894, 2.35%)
4 (894, 29.87%) (894, 40.04%) (894, 13.87%) (894, 2.68%) (894, 13.20%) (894, 2.35%)
6 (894, 29.75%) (894, 38.81%) (894, 12.75%) (894, 2.68%) (894, 13.20%) (894, 2.35%)
8 (892, 29.82%) (892, 36.77%) (892, 12.22%) (892, 2.69%) (892, 13.12%) (892, 2.24%)
As predictors to our model, we selected covariates that are typically monitored in the ICU and important for determining sepsis intervention strategies, as well as potential confounders. The covariates we used were similar to that of Li et al. Li et al. (2021), encompassing, but not limited to, basic demographic information, an Elixhauser comorbidity score, a SOFA score, laboratory values and vital signs, and urine output Li et al. (2021). A comprehensive list is provided in Tables 9 and 10. The demographics, comorbidities, and pre-ICU fluids were unmodeled and regarded as static while the remaining variables were modeled and regarded as dynamic (time-varying).
@@ E.1 Settings in Cancer Growth Experiments
Number of Layers (Representation) 1, 2
Hidden Dimension (Representation) 8, 16, 32, 64
@@ E.3 Settings in MIMIC Data
[…]
@@
Number of Layers (Continuous) 2, 3
Hidden Dimension (Categorical) 64
Hidden Dimension (Continuous) 64, 128
Batch Size 16
Weight Decay 1e-6
G-Net Learning Rate 0.0001
Number of Layers 2, 3
Hidden Dimension 64, 128
Batch Size 16
Weight Decay 1e-5
G-Transformer Learning Rate 0.0001
@@ E.4 Settings for Causal Transformer
Different from the Causal Transformer paper, we did not generate different counterfactual trajectories on the Cancer Growth dataset at each time step using Single sliding treatment or Random trajectories. For each patient in the test set, we inserted our counterfactual treatment at the last 5 time steps and applied this counterfactual treatment in each subsequent time step. In addition, for multi-step time prediction with Causal Transformer, we predicted the next 4 time steps and $\tau=5$ in this dataset. In both the CVSim dataset and the MIMIC dataset, since the model setup does not allow access to the actual treatment, we modified parts of the Causal Transformer’s code to enable the model to use the predicted treatment as input for predicting the outcomes of the next time step.

### v4

@@ 5.5 Counterfactual Experiments
We used G-Transformer to simulate covariate trajectories for patients in the test dataset under one conservative fluids strategies (with 5L fluid cap) and a fluid liberal strategy (with no cap). Selected covariates are presented in Figure 5, allowing for comparison of the two counterfactual regimes. Population-level average trajectories were calculated from the test set (N=894). Counterfactual strategies are applied starting at time step 1, conditioned on observations up until the first hour in the ICU. For each patient, we perform 100 Monte Carlo simulations. The plotted values are averaged across simulated trajectories for all test patients.
@@
Figure 5: Illustrative examples of population-level trajectories of selected covariates under counterfactual fluids strategies. G-Transformer’s predictions for selected covariates under various counterfactual fluids strategies, including one fluids conservative (purple) strategy with 5L fluids cap, and a fluids liberal (green) strategy without fluid cap.
We provide illustrative examples of patient trajectories at a population-level to demonstrate G-Transformer’s capability to generate clinically meaningful counterfactual predictions. The trends depicted in Figure 5 generally align with our physiological and clinical expectations. Under fluid liberal strategies, the bolus volume and blood pressure exhibit higher values, following the expected trends at a population level. Conversely, the higher levels in creatinine under the fluids conservative regime also align with expected outcomes.
Readers should exercise caution when interpreting these plots in Figure 5 as indicators of treatment effects, as the number of patients (or trajectories) contributing to the average may not remain constant throughout the entire 24-hour period. Simulated trajectories may end prematurely within the 24-hour period due to death or discharge.
@@ 6 Discussion and Conclusion
In this paper, we present G-Transformer, a Transformer-based framework supporting g-computation, to estimate clinical outcomes under counterfactual treatment strategies using both synthetic datasets and a real-world sepsis patient cohort. Our experiments using Cancer Growth and CVSim data, demonstrated that G-Transformer achieved better overall performance compared to other state-of-the-art methods. Using a real-world sepsis dataset, we demonstrated G-Transformer’s potential clinical utility in forecasting covariate trajectories under alternative counterfactual fluid limiting regimes. Although we focus on applications in healthcare in this study, the G-Transformer architecture can be applied to other sequential decision making tasks that involve dynamic and time-varying interventions.
@@
7 Acknowledgements
The authors are grateful for Professor Roger Mark for his insightful comments, and Stephanie Hu for her contribution in developing an earlier version of the sepsis cohort. L Lehman was in part funded by MIT-IBM Watson AI Lab, and NIH grant R01EB030362.
@@ A.1 Alternative G-Transformer Architecture Experiment
Our proposed G-Transformer training process differs from the traditional Transformer, in the way it is iteratively trained on time-series data. Conditioning on observations from the first time step, we treat each time step starting from the second one as an independent subsequence, and use this subsequence to predict the next time step. Then, we concatenated the results of the last time step of each sub-sequence to form a complete prediction sequence for optimization. In contrast, traditional Transformers directly use a sequence-to-sequence framework, shifting the training sequence to obtain the sequence that needs to be predicted. The reason for doing this is to meet the needs of g-computation and dynamic prediction. While g-computation calculates probability distributions based on the sequence leading up to the time step being predicted, the Transformer is a typical end-to-end model. We compared the performance of models implemented with these two training methods on the CVSim dataset, and the results are shown in Table 5.
@@
A.2 Individual-Level RMSE Over Time
@@
Appendix B MIMIC Experiments
@@ Appendix C CVSim Data Generation
We follow the same procedure as in Li et al. (2021) in generating the CVSim dataset. Briefly, a CVSim 6-compartment circulatory model takes as inputs 28 variables that together govern a hemodynamic system. It then deterministically simulates forward in time a set of 25 output variables according to a collection of differential equations (parameterized by the input variables) modeling hemodynamics. Important variables in CVSim include arterial pressure (AP), central venous pressure (CVP), total blood volume (TBV), and total peripheral resistance (TPR). In real patients, physicians observe AP and CVP and seek to keep them above a clinically safe threshold. They do this by intervening on TBV (through fluid administration) and TPR (through vasopressors).
We defined simulated treatment interventions that were designed to mimic the impact of fluids and vasopressors. These simulated interventions alter the natural course of the simulation by increasing either TBV (in the case of the simulated fluids intervention) or TPR (in the case of the simulated vasopressor intervention). We generated patients by randomly initiating baseline inputs (which we hid from our models to make this a stochastic modeling problem) within plausible physiologic ranges, then using CVSim to simulate covariates forward, intervening according to the relevant treatment strategy at each timestep.
@@
Appendix D MIMIC Data and Cohort
@@
k Edema MV Diuretics Dialysis
2 (894, 29.98%) (894, 40.94%) (894, 15.10%) (894, 2.68%)
6 (894, 29.75%) (894, 38.81%) (894, 12.75%) (894, 2.68%)
As predictors to our model, we selected covariates that are typically monitored in the ICU and important for determining sepsis intervention strategies, as well as potential confounders. The covariates we used were similar to that of Li et al. Li et al. (2021), encompassing, but not limited to, basic demographic information, an Elixhauser comorbidity score, a SOFA score, laboratory values and vital signs, and urine output Li et al. (2021). A comprehensive list is provided in Tables 8 and 9. The demographics, comorbidities, and pre-ICU fluids were regarded as static while the remaining variables were modeled and regarded as dynamic (time-varying).
@@ E.1 Settings in Cancer Growth Experiments
Number of Layers 1, 2
Hidden Dimension (Continuous) 8, 16, 32, 64
@@ E.3 Settings in MIMIC Data
Hyperparameter optimization on MIMIC data was performed by searching over the hyperparameter space shown in Table 12 with optimal parameters starred. All models were trained using the Adam optimizer with early stopping with patience of 10 epochs for a maximum of 50 epochs. For each covariate, the corresponding Transformer encoder in G-Transformer and LSTM in G-Net were optimized through a grid search of the following hyperparameters. The experiments were performed on NVIDIA Tesla T4 GPU with 8 vCPUs + 15GB of dedicated memory.
@@
Number of Layers 2
Hidden Dimension 64
Batch Size 32
Weight Decay 0.001
G-Net Learning Rate 0.001, 0.0001, 0.00001
Number of Layers 2
Hidden Dimension 64
Batch Size 32
Weight Decay 0.001
G-Transformer Learning Rate 0.001, 0.0001, 0.00001
@@ E.4 Settings for Causal Transformer
Different from the Causal Transformer paper, we did not generate different counterfactual trajectories on the Cancer Growth dataset at each time step using Single sliding treatment or Random trajectories. For each patient in the test set, we inserted our counterfactual treatment at the last 5 time steps and applied this counterfactual treatment in each subsequent time step. In addition, for multi-step time prediction with Causal Transformer, we predicted the next 4 time steps and $\tau=5$ in this dataset. In both the CVSim dataset and the MIMIC dataset, since the experiment setting does not allow access to the actual treatment, we modified parts of the Causal Transformer’s code to enable the model to use the predicted treatment as input for predicting the outcomes of the next time step. Specifically, we set the $\tau$ of Causal Transformer to 1, and repeat the entire prediction process $n$ times. Between each prediction step, we use a function to calculate whether the treatment is needed. The obtained treatment and the outcome predicted are then added as inputs to supplement the end of the input queue.